OpenStudy (anonymous):

Solve for x: |x| − 8 = −5 x = −13 and x = −3 x = 3 and x = −3 x = 3 and x = 13 No solution

OpenStudy (anonymous):

x - 8 = -5 x - 8 + 8 = -5 + 8 x = -5 + 8 x = 3

OpenStudy (anonymous):

there is 2 answers. you gave me one